OpenStudy (anonymous):

Given the ordered pair (3, y), what value of y would make the ordered pair a solution of the equation 4x − 2y = 24?

OpenStudy (anonymous):

A. 12 B. 6 C. -6 D. -8

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Find the answer by putting x=3, then find y.

OpenStudy (anonymous):

4x − 2y = 24 x=3 So we just put the x there 4*3-2y=24 12-2y=24 -2y=24-12 -2y=12 y=-6